# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

### v4

@@ 1 Introduction
Supplementing prior instances such as Santa Fe processes introduced in [1, 2], the aim of the present paper is to construct further examples of stationary sources over a countable alphabet that obey a power-law growth of block entropy, called Hilberg‘s law [3]. The processes constructed in this paper are called multiperiodic processes. Multiperiodic processes are stationary ergodic (but not mixing) and enjoy the vanishing entropy rate—in contrast to Santa Fe processes [1, 2], which are non-ergodic and usually have a positive entropy rate. As it has been demonstrated in [4], Hilberg‘s law—deemed to hold for natural language [3]—implies the better known neural scaling law, which is a power-law decay of the cross entropy rate and empirically holds for large language models trained on internet-sized corpora of texts [5, 6, 7, 8]. We hope that the toy examples of Santa Fe and multiperiodic processes may be useful as baseline artificial data sources to be compared with natural texts in quantitative linguistics and machine learning.
We cast our contribution against the background of previous research.
Preliminaries.
@@
For a stationary process $(X_{t})_{t\in\mathbb{Z}}$, we may define two limits
@@
Limit $h$ is called entropy rate [9, 10], whereas $E$ is called excess entropy [11] or predictive information [12, 13, 14]. While entropy rate is a measure of process unpredictability, excess entropy is an intuitive measure of process memory. Although excess entropy only measures the memory capacity, without specifying how exactly the process future depends on the past, it can be given interesting interpretations. Some attempts were also made to generalize excess entropy to two-dimensional random fields [15, 16].
@@
Conditions $E<\infty$ and $E=\infty$ have been studied in probability theory and its applications for quite a time. The oldest systematic results seem to concern Gaussian processes. In particular, Grenander and Szegő [17, Section 5.5] presented an integral formula for excess entropy (in disguise). Finch [18] evaluated this formula for autoregressive moving average (ARMA) processes, which turned out to yield $E<\infty$. More generally, we have $E<\infty$ if and only if $\sum_{k=1}^{\infty}k\alpha_{k}^{2}<\infty$ and $\left|\alpha_{k}\right|<1$ for the partial autocorrelation $\alpha_{k}$ [19]. In particular, under a positive and continuous spectral density, Li [20] showed that $E<\infty$ if and only if the autocorrelation $\rho_{k}:=\operatorname{Corr}(X_{0};X_{k})$ satisfies $\sum_{k=1}^{\infty}k\rho_{k}^{2}<\infty$.
Let us proceed to the countable alphabet. It can be easily shown, by the data-processing inequality, that condition $E<\infty$ holds for hidden Markov processes with a finite number of states [11]. Mahoney, Ellison and Crutchfield [21, 22] developed a formula for excess entropy in terms of predictive and retrodictive $\epsilon$-machines, which are minimal unifilar hidden Markov representations of the process [23, 24]. Travers and Crutchfield [25] and Dębowski [26] exhibited examples of hidden Markov processes with a countably infinite number of hidden states that obey $E=\infty$ and the further discussed Hilberg‘s law (4). A few other older papers concerned also isolated examples of processes with $E=\infty$ over a finite alphabet. For instance, Bradley [27] constructed the first example of a mixing process having this property. Gramss [28] investigated a process which is formed by the frequencies of 0‘s and 1‘s in the rabbit sequence.
@@
where $\beta\in(0,1)$, holds for natural language. Although the first reception of Hilberg‘s law (4) came from physicists [30, 31, 13, 11, 24], there is a line of research that seeks to combine a rigorous information-theoretic approach to Hilberg‘s law with general linguistic insights [32, 1, 2, 33, 34, 35] such as Zipf‘s law in particular [36, 37]. The basic observation is that condition $E=\infty$ holds if the process is strongly non-ergodic in the sense that its shift-invariant algebra $\mathcal{I}$ is non-atomic. It is so because
@@
There may exist connections between the Santa Fe decomposition (6) and algorithmic information theory [44], see also [34, Chapter 1]. In the role of immutable knowledge, the random probability $Y:=\sum_{k=1}^{\infty}2^{-k}Z_{k}\in(0,1)$ resembles Chaitin‘s number $\Omega\in(0,1)$, also called the halting probability [45, 46]. The number $\Omega$ is algorithmically random [47] and encodes a large body of mathematical knowledge in a dense form. However, equating the values of probabilities $Y$ and $\Omega$ is preposterous for two reasons. The first of them, a formal one, is that $Y$ is a random variable, whereas $\Omega$ is a fixed number. The second reason, philosophically more important, is that $\Omega$ does not seem computable in a sufficiently strong sense to arise as an effect of natural processes [48]. If there are some links between the Santa Fe decomposition (6) and algorithmic information theory, they must be more subtle.
@@
where $\alpha>1$ [36, 37]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (4) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^1] We adopt the distinction of terms “type” and “token” borrowed from quantitative linguistics [49, 50]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
@@
The above described Santa Fe process is non-ergodic and takes values in a countably infinite alphabet. The reason for its non-ergodicity is that coin flips $(Z_{k})_{k\in\mathbb{N}}$ are measurable with respect to the shift-invariant algebra of process $(X_{t})_{t\in\mathbb{Z}}$. Thus each particular value of sequence $(Z_{k})_{k\in\mathbb{N}}$ defines a distinct ergodic component [1, 34]. However, replacing individual coin flips $Z_{k}$ with slowly evolving Markov chains $(Z_{kt})_{t\in\mathbb{Z}}$, we obtain processes which are (strongly) mixing hence ergodic [33]. Coding natural numbers as binary strings, we obtain processes over a finite alphabet [2, 35]. All these examples enjoy Hilberg‘s law (4) with a positive rate $h>0$.
@@ 2.1 Multiperiodic sequences
Multiperiodic sequences are an idea which we have originally generalized from work [58]. The authors only considered a fixed multiperiodic sequence with particular periods and seeds—to state a certain negative result in universal prediction. Isolated instances of multiperiodic sequences are noted in the on-line encyclopedia of integer sequences [59] and can be connected to other problems, such as Mancala-type games [60]. Distilling the general construction, especially considering the idea of random seeds that matters in the probabilistic context, seems our contribution.
@@ 2.2 Multiperiodic processes
Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [61, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
@@
for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [34, Section 4.1].
@@
Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [34, Theorem 4.10] or [62, Lemma 7.15]. ∎
@@ 3.3 Number of types
Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [61, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (52) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (53). ∎
@@ 5.2 Linear periods
To provide some visualization of these results, in Figure 1, we present the growth of the actual number of types $V_{t}$ for three multiperiodic sequences with periods $\pi_{k}=1+\left\lceil ck\right\rceil$ and seeds $\sigma_{k}=1$, where $c\in\left\{0.5,1.0,2.0\right\}$. We notice a good agreement with the theoretical prediction $\operatorname{\mathbf{E}}V_{t}\sim t^{c/(c+1)}$.

### v5

@@ 1 Introduction
Power-law regularities in the growth of block entropy, succinctly referred to as Hilberg‘s law [1], have long been conjectured to characterize natural language and related symbolic data [2, 3, 4, 5]. As it has been recently demonstrated in [6], Hilberg‘s law implies the better known neural scaling law, which is a power-law decay of the cross entropy rate of a large language model with respect to its training resources. The latter law empirically holds for internet-sized corpora of texts [7, 8, 9, 10]. From a theoretical perspective, these connections motivate the construction of explicit stochastic sources that obey Hilberg‘s law and can serve as analytically tractable reference models.
Earlier work introduced the Santa Fe and Oracle processes [11, 4, 12, 13] as some simple examples of stationary sources over a countable alphabet exhibiting power-law block entropy growth, albeit at the cost of non-ergodicity and a positive entropy rate. The present paper supplements these constructions by introducing a new class of processes, termed multiperiodic processes. These processes are stationary and ergodic, though not mixing, and they exhibit a vanishing entropy rate while still obeying Hilberg‘s law. We note that the construction of Santa Fe processes is quite simple, whereas that of multiperiodic processes is a somewhat more technical but still elementary.
We hope that both Santa Fe and multiperiodic processes may serve as useful toy models or baseline artificial data sources for quantitative comparisons with natural language in information theory, linguistics, and machine learning. The practical advantage of these models is that we can fully control the entropy rate and the Hilberg law exponent [14] of an easily simulated process. This aspect matters if we want to empirically check or theoretically prove the convergence rate of some entropy estimators, universal codes, or deep learning algorithms.
Preliminary concepts.
@@
For a stationary process $(X_{t})_{t\in\mathbb{Z}}$, we consider two limits
@@
Limit $h$ is called the entropy rate [15, 16], whereas $E$ is called the excess entropy [3] or the predictive information [17, 18, 19]. While entropy rate is a measure of process unpredictability, excess entropy is an intuitive measure of process memory. Although excess entropy only measures the memory capacity, without specifying how exactly the process future depends on the past, it can be given interesting interpretations. Some attempts were also made to generalize excess entropy to two-dimensional random fields [20, 21].
@@
Conditions $E<\infty$ and $E=\infty$ have been studied in probability theory and its applications for quite a time. The oldest systematic results seem to concern Gaussian processes. In particular, Grenander and Szegő [22, Section 5.5] presented an integral formula for excess entropy (in disguise). Finch [23] evaluated this formula for autoregressive moving average (ARMA) processes, which turned out to yield $E<\infty$. More generally, we have $E<\infty$ if and only if $\sum_{k=1}^{\infty}k\alpha_{k}^{2}<\infty$ and $\left|\alpha_{k}\right|<1$ for the partial autocorrelation $\alpha_{k}$ [24]. In particular, under a positive and continuous spectral density, Li [25] showed that $E<\infty$ if and only if the autocorrelation $\rho_{k}:=\operatorname{Corr}(X_{0};X_{k})$ satisfies $\sum_{k=1}^{\infty}k\rho_{k}^{2}<\infty$.
Let us proceed to the countable alphabet. It can be easily shown, by the data-processing inequality, that condition $E<\infty$ holds for hidden Markov processes with a finite number of states [3]. Mahoney, Ellison, and Crutchfield [26, 27] developed a formula for excess entropy in terms of predictive and retrodictive $\epsilon$-machines, which are minimal unifilar hidden Markov representations of the process [28, 29]. Travers and Crutchfield [30] and Dębowski [31] exhibited examples of hidden Markov processes with a countably infinite number of hidden states that obey $E=\infty$ and the further discussed Hilberg‘s law (4). A few other older papers concerned also isolated examples of processes with $E=\infty$ over a finite alphabet. For instance, Bradley [32] constructed the first example of a mixing process having this property. Gramss [33] investigated a process which is formed by the frequencies of 0‘s and 1‘s in the rabbit sequence.
@@
where $\beta\in(0,1)$, holds for natural language. Although the first reception of Hilberg‘s law (4) came from physicists [2, 35, 18, 3, 29], there is a line of research that seeks to combine a rigorous information-theoretic approach to Hilberg‘s law with general linguistic insights [36, 11, 4, 37, 14, 38, 13] such as Zipf‘s law in particular [39, 40]. The basic observation is that condition $E=\infty$ holds if the process is strongly non-ergodic in the sense that its shift-invariant algebra $\mathcal{I}$ is non-atomic. It is so because
@@
There may exist connections between the Santa Fe decomposition (6) and algorithmic information theory [46], see also [38, Chapter 1]. In the role of immutable knowledge, the random probability $Y:=\sum_{k=1}^{\infty}2^{-k}Z_{k}\in(0,1)$ resembles Chaitin‘s number $\Omega\in(0,1)$, also called the halting probability [47, 48]. The number $\Omega$ is algorithmically random [49] and encodes a large body of mathematical knowledge in a dense form. However, equating the values of probabilities $Y$ and $\Omega$ is preposterous for two reasons. The first of them, a formal one, is that $Y$ is a random variable, whereas $\Omega$ is a fixed number. The second reason, philosophically more important, is that $\Omega$ does not seem computable in a sufficiently strong sense to arise as an effect of natural processes [50]. If there are some links between the Santa Fe decomposition (6) and algorithmic information theory, they must be more subtle.
@@
where $\alpha>1$ [39, 40]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (4) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^1] We adopt the distinction of terms “type” and “token” borrowed from quantitative linguistics: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
@@
The above described Santa Fe process is non-ergodic and takes values in a countably infinite alphabet. The reason for its non-ergodicity is that coin flips $(Z_{k})_{k\in\mathbb{N}}$ are measurable with respect to the shift-invariant algebra of process $(X_{t})_{t\in\mathbb{Z}}$. Thus each particular value of sequence $(Z_{k})_{k\in\mathbb{N}}$ defines a distinct ergodic component [11, 38]. However, replacing individual coin flips $Z_{k}$ with slowly evolving Markov chains $(Z_{kt})_{t\in\mathbb{Z}}$, we obtain processes which are (strongly) mixing hence ergodic [37]. Coding natural numbers as binary strings, we obtain processes over a finite alphabet, such as Oracle processes [4, 56, 13]. All these examples enjoy Hilberg‘s law (4) with a positive rate $h>0$.
@@ 2.1 Multiperiodic sequences
Multiperiodic sequences are an idea which we have originally generalized from work [61]. These authors only considered a fixed multiperiodic sequence with particular periods and seeds—to state a certain negative result in universal prediction. Isolated instances of multiperiodic sequences are noted in the on-line encyclopedia of integer sequences [62] and can be connected to other problems, such as Mancala-type games [63]. Distilling the general construction, especially considering the idea of random seeds that matters in the probabilistic context, seems our contribution.
@@ 2.2 Multiperiodic processes
Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [64, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
@@
for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [38, Section 4.1].
@@
Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [38, Theorem 4.10] or [65, Lemma 7.15]. ∎
@@ 3.3 Number of types
Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [64, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (52) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (53). ∎
@@ 5.2 Linear periods
To provide some visualization of these results, in Figure 1, we present the growth of the actual number of types $V_{t}$ for three multiperiodic sequences with periods $\pi_{k}=1+\left\lceil ck\right\rceil$ and seeds $\sigma_{k}=1$, where $c\in\left\{0.25,0.5,1.0,2.0,4.0\right\}$. We notice a good agreement with the theoretical prediction $\operatorname{\mathbf{E}}V_{t}\sim t^{c/(c+1)}$.
@@ 6 Conclusion
AI application
The style of the first three paragraphs in the Introduction has been improved by means of ChatGPT. The suggestions by the model have been post-edited.